# ILRe: Intermediate Layer Retrieval for Context Compression in Causal Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated success across many benchmarks. However, they still exhibit limitations in long-context scenarios, primarily due to their short effective context length, quadratic computational complexity, and high memory overhead when processing lengthy inputs. To mitigate these issues, we introduce a novel context compression pipeline, called Intermediate Layer Retrieval (ILRe), which determines one intermediate decoder layer offline, encodes context by streaming chunked prefill only up to that layer, and recalls tokens by the attention scores between the input query and full key cache in that specified layer. In particular, we propose a multi-pooling kernels allocating strategy in the token recalling process to maintain the completeness of semantics. Our approach not only reduces the prefilling complexity from $O(L^{2})$ to $O(L)$, but also achieves performance comparable to or better than the full context in the long context scenarios. Without additional post training or operator development, ILRe can process a single $1M$ tokens request in less than half a minute (speedup $\approx 180\times$) and scores RULER-$1M$ benchmark of $\approx 79.8$ with model Llama-3.1-UltraLong-8B-1M-Instruct on a Huawei Ascend 910B NPU.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success across a wide range of context processing tasks, including document retrieval (Laban et al., 2023), code generation (Gu, 2023), multimodal applications (Sapkota et al., 2025), and agent systems  (Guo et al., 2024; Luo et al., 2025). Especially in the long context scenarios, many works (Liu et al., 2025) are contributed to explore the limits of abilities of LLMs. Meanwhile, mainstream LLMs are increasingly supporting longer context lengths (Yang et al., 2025; Xu et al., 2025; MiniMax et al., 2025).
@@
Table 1: RULER benchmark performance of different methods with Llama-3.1-UltraLong-8B-1M-Instruct. Except the FullContext, the token budget is set to be 4K. The best scores are in bold faces.
@@
Figure 1: Time To First Token(TTFT) of Intermediate Layer Retrieval (ILRe) with Llama-3.1-8B on a Huawei Ascend 910B NPU. The red text notate the speedup of our approach.
Recently, many works have been contributed to mitigate the above issues. SnapKV (Li et al., 2024b) and Quest (Tang et al., 2024) proposed using the attention scores of an observation window to select salient tokens from the KV cache to reduce the complexity of decoding phase. DuoAttention (Xiao et al., 2025) introduced a streaming chunked prefill to encode incoming tokens within streaming heads in the prefilling phase. However, integrating these techniques into a unified, highly accurate, and efficient pipeline remains an open challenge. Meanwhile, our empirical observations reveal that different decoder layers exhibit varying recall rates of deserved content retrieved by SnapKV-like operation in which we use maximums of attention weights to unify focused content. As illustrated in the “One Kernel” series in Figure 2, some intermediate layers have high recall rates which also exhibit degradation when the length of key increases (same series in Figure 3).
Figure 2: Recall of layers in model Llama-3.1-UltraLong-8B-1M-Instruct over 128K context. The context is prefilled by streaming chunked prefill and the scores are averaged over multiple depths and key lengths.
Figure 3: Recall of different Key Lengths in model Llama-3.1-UltraLong-8B-1M-Instruct over 128K context. The context is prefilled by streaming chunked prefill and the scores are averaged over multiple depths.
Inspired by the works and observations mentioned above, we propose a novel context compression method, Intermediate Layer Retreival(ILRe), as depicted in Figure 4. Our method processes the long full context with partial activated model parameters and streaming chunked prefill to reduce the prefilling complexity. In the layer determined by a simple retrieval task, it collects all the past key states of each chunk and the query states of question so that it can perform a SnapKV-like operation to get full attention weights for query-aware global analysis. Specifically, we propose a strategy that evenly gathers retrieval tokens from multiple combinations of max- and avg-pooling to enhance the capability of retrieving long contents (“Multi Kernels” series in Figure 2 and  3).
@@
We evaluate our approach with various benchmarks, including RULER and LongBench (Bai et al., 2023), to demonstrate its effectiveness. The results show that our approach achieves remarkable performance across all benchmarks over other compression methods.
@@ 2.1 Internal Memory Compression
Autoregressive LLMs store key–value (KV) pairs during decoding to avoid recomputation, but this cache grows linearly with sequence length, creating a significant memory bottleneck. Several approaches have been proposed to mitigate this issue by selectively retaining or compressing the stored representations. H2O (Zhang et al., 2024) preserves only “heavy hitter” tokens that dominate attention scores, while FastGen (Ge et al., 2024) applies runtime profiling to dynamically choose compression policies. SnapKV (Li et al., 2024c) pools attention scores across prompt tokens to identify and retain informative clusters. Other works identify attention heads specialized for long-range recall and introduce head-specific cache compression (Wu et al., 2024; Fu et al., 2024b; Liu et al., 2024a; Jiang et al., 2024a).
@@ 3.1 Preliminaries
where $PE$ stands for the position encoding. The attention is calculated as:
@@ 3.2 Observations of Retrieval Tasks
As demonstrated by SnapKV, the critical components of the KV cache for a specific query can be identified through the attention between the context and the query, which is calculated as:
@@
Here, we assume that the batch size is $1$, that is, $A_{h,l}^{q,C}$ is a 3D matrix with head, query, and context dimensions. Let $L_{q}$ denote the size of the query $q$; typically, $L_{q}$ is much smaller than the length $L$ of the context $C$. In the process of context compression, we need to determine which input IDs of context (rather than KV pairs) should be retained for re-inference. To achieve this, we must reduce the attention from 3D to a 1D matrix with only the context dimension left. For that, we only keep the maximums of attention across the head and query dimensions in one specific layer $l_{R}$:
@@
In the simple retrieval task defined in Appendix A.1, we calculate the recall rate of each layer across varying key depths and lengths, as presented in Figure 2 and Figure 3. We observe that certain layers equipped with only a single average kernel can achieve a $100\%$ recall rate in short-key tasks but decline sharply as the key size increases. In contrast, our multi-kernel approach maintains consistent performance across different key sizes. Furthermore, frontier layers exhibit higher recall rates than later layers, resulting in the optimal $l_{R}$ typically being smaller than half of $N_{L}$ and thus yielding much lower computational complexity. We define the optimal layer $l_{R}$ as the layer with the smallest index value among those exhibiting the highest recall rates. For example, in the model Llama-3.1-Ultralong-8B-1M-Instruct (Xu et al., 2025), we identify the optimal $l_{R}=3$ versus the total number of layers $N_{L}=32$. Meanwhile, we count the frequency of occurrence of maxima arguments of attention heads in Figure 5, which demonstrates that attention heads are not fixed in their behavior.
Figure 5: Occurrence Ratios of max head indices in one $128K$ Retrieval Task of model Llama-3.1-Ultralong-8B-1M-Instruct.
@@ 3.3 Intermediate Layer Retrieval
  • Determine $l_{R}$ offline as discussed above.
  • Load LLM parameters up to the layer $l_{R}$. Only these parameters will be used for the context compression.
  • Prepare a long context and query part. Usually, the query part consists of a question and answer prefix.
@@
  • In the retrieval layer $l_{R}$, calculate the key states for the context and the query states for the query part, and then skip the rest of the calculations. Pre-allocate memory with context length $L$ for key states caching $K^{C}_{h,l_{R}}$ and fill the cache chunk by chunk.
@@
  • Calculate $A^{C}$ with the cached key states and the query states.
@@ 3.4 Complexity Analysis
We only discuss the attention complexity of context compression, as the standard inference of a compressed prompt is constrained by a fixed budget $B$.
@@ 4 Experiments
Single Doc. QA Multi Doc. QA Summarization Fewshot Learning Synthetic Code
$B$ Method
@@
Table 3: Longbench results of model Llama-3.1-8B-Instruct. The best and second best scores are boldfaced and underlined respectively.
Baseline Methods. For comparison, except the original prompt with full context, we select other methods similar to or used in the components of our approaches: StreamingLLM (Xiao et al., 2023), SnapKV (Li et al., 2024b), and Retrieval-Augmented Generation (RAG) (Gao et al., 2024; Gupta et al., 2024) with Qwen3-Embedding-0.6B (Zhang et al., 2025) as the backbone.
Benchmarks and LLMs. In this work, we utilize two benchmarks: LongBench and RULER. For RULER, we set the token budget to $4K$. Given that LongBench contexts are relatively short (with an average length of fewer than $16K$ tokens), we only test the Llama-3.1-8B-Instruct model (Grattafiori et al., 2024)—which claims a $128K$ context capacity—setting the token budget to $1K$ and $2K$. We also test an additional model family: Qwen2.5-7B-Instruct (Qwen et al., 2025). Similar to Figure 2 and  3, the results of Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct in the Retrieval Task are present in Figure 6 and  7 respectively.
Figure 6: Results of model Llama-3.1-8B-Instruct with $128K$ context in Retrieval Trask.
Figure 7: Results of model Qwen2.5-7B-Instruct with $32K$ context in Retrieval Trask.
Default Settings of ILRe. By default, we set $S=4$, $W=512$, and the chunk size of prefilling is 10241024(the first chunk with additional $S$). For the pooling, we use $3$ max pooling kernels with sizes $2,4,8$ and $16$ average pooling kernels with sizes from $1$ to $16$ as mentioned in Section 3.2.
4.1 RULER
4.2 LongBench
@@
We highlight two key observations. First, ILRe achieves improvements over the full context at 2K token budget. ILRe with full KV retention attains the best overall performance, outperforming SnapKV and StreamingLLM in most tasks, with particularly strong gains in retrieval-heavy tasks such as QA and summarization. Even with various layers (e.g. $l_{R}=3,6,10,11$), ILRe remains competitive, demonstrating that intermediate layer retrieval can recover useful long-context information with fewer stored tokens. Second, at 1K context length, ILRe continues to scale favorably, showing robustness under tighter memory budgets.
@@ 4.3 Ablations
In this part, we will change the settings or modify the components of the standard ILRe to prove its robustness.
@@
Table 4: RULER benchmark performance of different prefilling methods with Llama-3.1-UltraLong-8B-1M-Instruct.
Prefilling Method. To prove the stability of the streaming chunked prefill in our framework, we compare it with using the full KV in layers $\{1,2,...,l_{R}-1\}$ by setting $W=\infty$. As demonstrated in Table 4, the results of RULER benchmark of the streaming chunked prefill are comparable to those of full KV. The results with “FullKV” in Table 3 further reinforce this finding.
@@
Table 5: RULER benchmark performance of different pooling settings with Llama-3.1-UltraLong-8B-1M-Instruct. The best of average scores are in bold faces.
Multi Kernels. We define three different configurations of the pooling method from the default setups:
  • Less-Max-Kernels uses only one max pooling with kernel size $4$.
  • Less-Avg-Kernels reduces the number of average pooling to nine with sizes ranging from $1$ to $9$.
  • One-Max-Avg only has one max pooling and one average pooling with size $4$ and $5$ respectively.
@@
8K 16K 32K
FullContext 93.1 91.4 84.7
ILRe with $l_{R}=6$ 87.2 85.1 69.6
ILRe with $l_{R}=5$ 93.6 91.9 87.9
Table 6: RULER benchmark performance of ILRe with Qwen2.5-7B-Instruct. The best scores are in bold faces.
Other Models. For the model Qwen2.5-7B-Instruct, we identify the optimal $l_{R}=5$ and limit testing to $32K$ in Table 6, consistent with its claimed capabilities. This result demonstrates that our approach is valid across other model families.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated success across many benchmarks. However, they still exhibit limitations in long-context scenarios, primarily due to their short effective context length, quadratic computational complexity, and high memory overhead when processing lengthy inputs. To mitigate these issues, we introduce a novel context compression pipeline, called Intermediate Layer Retrieval (ILRe), which determines one intermediate decoder layer offline, encodes context by streaming chunked prefill only up to that layer, and recalls tokens by the attention scores between the input query and full key cache in that specified layer. In particular, we propose a multi-pooling kernels allocating strategy in the token recalling process to maintain the completeness of semantics. Our approach not only reduces the prefilling complexity from $O(L^{2})$ to $O(L)$ and trims the memory footprint to a few tenths of that required for the full context, but also delivers performance comparable to or superior to the full-context setup in long-context scenarios. Without additional post training or operator development, ILRe can process a single $1M$ tokens request in less than half a minute (speedup $\approx 180\times$) and scores RULER-$1M$ benchmark of $\approx 79.8$ with model Llama-3.1-UltraLong-8B-1M-Instruct on a Huawei Ascend 910B NPU.
@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable success across a wide range of context processing tasks, including document retrieval (Laban et al., 2023), code generation (Gu, 2023), multimodal applications (Sapkota et al., 2025), and agent systems  (Guo et al., 2024; Luo et al., 2025). Especially in the long context scenarios, many works (Liu et al., 2025) are contributed to explore the limits of abilities of LLMs. Meanwhile, mainstream LLMs are increasingly supporting longer context lengths (Yang et al., 2025b; Xu et al., 2025; MiniMax et al., 2025).
Table 1: RULER benchmark performance of different methods with Llama-3.1-UltraLong-8B-1M-Instruct. Except the FullContext, the token budget is set to be 4K. Dual Chunk Attention (DCA) is optional for ILRe. The best scores are in bold faces.
@@
ILRe w/ DCA (Ours) 86.8 83.5 83.5 84.3 84.0
@@
Figure 1: Time To First Token(TTFT) of Intermediate Layer Retrieval (ILRe) with Llama-3.1-8B on a Huawei Ascend 910B NPU. The red text notates the speedup of our approach.
Recently, many works have been contributed to mitigate the above issues. SnapKV (Li et al., 2024b) and Quest (Tang et al., 2024) proposed using the attention scores of an observation window to select salient tokens from the KV cache to reduce the complexity of decoding phase. DuoAttention (Xiao et al., 2025) introduced a streaming chunked prefill to encode incoming tokens within streaming heads in the prefilling phase. However, integrating these techniques into a unified, highly accurate, and efficient pipeline remains an open challenge. Meanwhile, our empirical observations reveal that different decoder layers exhibit varying recall rates of deserved content retrieved by SnapKV-like operation in which we use maximums of attention weights to unify focused content. As illustrated in the “One Kernel” series in Figure 2(a), some intermediate layers have high recall rates which also exhibit degradation when the length of key increases (same series in Figure 2(b)).
(a) Different layers.
(b) Different depths.
Figure 2: Retrieval Task results of model Llama-3.1-UltraLong-8B-1M-Instruct over 128K context. The context is prefilled by streaming chunked prefill.
Inspired by the works and observations mentioned above, we propose a novel context compression method, Intermediate Layer Retreival (ILRe), as depicted in Figure 3. Our method processes the long full context with partial activated model parameters and streaming chunked prefill to reduce the prefilling complexity. In the layer determined by a simple retrieval task, it collects all the past key states of each chunk and the query states of question so that it can perform a SnapKV-like operation to get full attention weights for query-aware global analysis. Specifically, we propose a strategy that evenly gathers retrieval tokens from multiple combinations of max- and avg-pooling to enhance the capability of retrieving long contents (“Multi Kernels” series in Figure 2).
@@
We evaluate our approach with two benchmarks, RULER and LongBench (Bai et al., 2023), to demonstrate its effectiveness. The results show that our approach achieves remarkable performance across all benchmarks over other compression methods.
@@ 2.1 Internal Memory Compression
Autoregressive LLMs store key–value (KV) pairs during decoding to avoid recomputation, but this cache grows linearly with sequence length, creating a significant memory bottleneck. Several approaches have been proposed to mitigate this issue by selectively retaining or compressing the stored representations. H2O (Zhang et al., 2024) preserves only “heavy hitter” tokens that dominate attention scores, while FastGen (Ge et al., 2024) applies runtime profiling to dynamically choose compression policies. SnapKV (Li et al., 2024c) pools attention scores across prompt tokens to identify and retain informative clusters. Other works identify attention heads specialized for long-range recall and introduce head-specific cache compression (Wu et al., 2024; Fu et al., 2024b; Liu et al., 2024a; Jiang et al., 2024a). Instead of compressing KV states, another line of research explores how to reuse them more effectively. CacheBlend, for example, integrates cached KV segments from earlier prompts into the current decoding process to save recomputation (Yao et al., 2025). Along this line, Block-attention (Ma et al., 2025) introduces a blockwise mechanism that divides retrieved documents into discrete blocks. This design allows previously seen blocks to reuse their KV states, reducing inference overhead in long-context scenarios like retrieval-augmented generation (RAG).
@@ 2.2 Prompt-Level Context Reduction
2.3 Context Length Extrapolation
A key challenge in deploying long-context models is their ability to generalize beyond their training sequence lengths(Zhao et al., 2024). Recent work has explored various approaches to enable length extrapolation, including positional encoding such as RoPE scaling(Su et al., 2023), YaRN(Peng et al., 2024). Other methods focus on architectural changes like sparse attention patterns(Lou et al., 2024) to extend context capabilities. While these approaches address the fundamental challenge of length generalization, they often require significant computational overhead or model retraining. Rather than scaling positions, Dual Chunk Attention (DCA)(An et al., 2024a) partitions long sequences into manageable chunks where all relative distances remain within the original training bounds, enabling models like LLaMA2-70B to handle contexts exceeding 100K tokens.
@@ 3.1 Preliminaries
where $PE$ stands for the positional encoding. The attention is calculated as:
@@
Rotary Positional Encodings (RoPE)(Sun et al., 2023) is currently the standard choice for $PE$ in LLMs. It’s formalized as:
$PE(\mathbf{X})=\mathbf{R}(m)\mathbf{X}$ (3)
where $m$ is the position index and $\mathbf{R}(m)$ is a rotation matrix.
@@ 3.2 Observations of Retrieval Tasks
As demonstrated by SnapKV, the critical components of the KV cache for a specific query $q$ can be identified through the attention between the context and the query, which is calculated as:
@@
Here, we assume that the batch size is $1$, that is, $A_{h,l}^{q,C}$ is a 3D matrix in the layer $l$ with head, query, and context dimensions. Let $L_{q}$ denote the size of the query $q$; typically, $L_{q}$ is much smaller than the length $L$ of the context $C$. In the process of context compression, we need to determine which input IDs of context (rather than KV pairs) should be retained for inference. To achieve this, we must reduce the attention from 3D to a 1D matrix with only the context dimension left. For that, we only keep the maximums of attention across the head and query dimensions in one specific layer $l_{R}$:
@@
In the simple retrieval task defined in Appendix A.1, we calculate the recall rate of each layer across varying key depths and lengths, as presented in Figure 2. We observe that certain layers equipped with only a single average kernel can achieve a $100\%$ recall rate in short-key tasks but decline sharply as the key size increases. In contrast, our multi-kernel approach maintains consistent performance across different key sizes. Furthermore, frontier layers exhibit higher recall rates than later layers, resulting in the optimal $l_{R}$ typically being smaller than half of $N_{L}$ and thus yielding much lower computational complexity. We define the optimal layer $l_{R}$ as the layer with the smallest index value among those exhibiting the highest recall rates. For example, in the model Llama-3.1-Ultralong-8B-1M-Instruct (Xu et al., 2025), we identify the optimal $l_{R}=3$ versus the total number of layers $N_{L}=32$. Meanwhile, we count the frequency of occurrence of maxima arguments of attention heads in Appedix A.2, which demonstrates that attention heads are not fixed in their behavior.
@@ 3.3 Intermediate Layer Retrieval
  • Determine $l_{R}$ offline as discussed above. Load LLM parameters up to the layer $l_{R}$.
  • Prepare a long context and short query part. Usually, the query part consists of a question and an answer prefix. Split the context $C$ into chunks $\{C_{i}\}_{i=1}^{N_{C}}$.
@@
  • In the retrieval layer $l_{R}$, calculate the key states for the context, and then skip the rest of the calculations. Pre-allocate memory with context length $L$ for key states caching $K^{C}_{h,l_{R}}$ and fill the cache with chunked key states $\{K^{C_{i}}_{h,l_{R}}\}_{i=1}^{N_{C}}$.
@@
  • Calculate $A^{C}$ with the cached key states and the query states in the layer $l_{R}$.
@@
Dual Chunk Attention (DCA)(An et al., 2024b). As proposed in DCA, the relative positional encoding with the bound distance can effectively extrapolate the length capacity of LLMs. We adapt DCA to ILRe in the following way. For the sink parts $K^{S}_{l,h}$ in layers $\{1,2,...,l_{R}-1\}$, we re-RoPE them by the chunk size as the distance $D_{C}$ in each chunk prefilling since the third chunk. That is
$K^{S,i}_{l,h}=\mathbf{R}^{i-2}(D_{C})K^{S}_{l,h}\text{ for }l\in\{1,2,...,l_{R}-1\}\text{ and }i>2\text{ only}.$ (6)
For the key states in the layer $l_{R}$, we re-RoPE them by the following equation before recalling tokens:
$\tilde{K}^{C_{i}}_{h,l_{R}}=\mathbf{R}(D_{i})K^{C_{i}}_{h,l_{R}}\text{ for }i<N_{C}-1\text{ only}.$ (7)
where $D_{i}$ is the distance between chunk $i$ and chunk $N_{C}-1$.
We select the chunk $N_{C}-1$ as the positional base in considering the last chunk possible shorter than others. So, the distance between all chunked key states and the query state will be bounded in the range of two times the chunk size.
@@ 3.4 Complexity Analysis
We only discuss the attention complexity in the context compression, as the standard inference of a compressed prompt is constrained by the fixed budget $B$.
@@ 4 Experiments
Baseline Methods. For comparison, except the original prompt with full context, we select other methods similar to or used in the components of our approaches: StreamingLLM (Xiao et al., 2023), SnapKV (Li et al., 2024b), and Retrieval-Augmented Generation (RAG) (Gao et al., 2024; Gupta et al., 2024) with Qwen3-Embedding-0.6B (Zhang et al., 2025) as the backbone. RAG uses $256$ tokens as the text chunk size.
Benchmarks and LLMs. In this work, we use two benchmarks: LongBench and RULER. For details of setups in these benchmarks, see Appendix A.3.
Default Settings of ILRe. By default, we set $S=4$, $W=512$, and the chunk size of prefilling is 10241024(the first chunk with additional $S$). For the pooling, we use $3$ max pooling kernels with sizes $2,4,8$ and $16$ average pooling kernels with sizes from $1$ to $16$ as mentioned in Section 3.2. DCA is off unless it is specified.
Specific Terms. The term “FullKV” refers to $W=\infty$ and “FullContext” for whole-context inference with full KV cache retention.
4.1 RULER
4.2 LongBench
Table 3: Longbench results of model Llama-3.1-8B-Instruct and Qwen2.5-7B-Instruct. The best and second best scores in each group are boldfaced and underlined respectively.
Single Doc. QA Multi Doc. QA Summarization Few-shot Learning Synthetic Code
Model $B$ Method
@@ 4.2 LongBench
Qwen2.5-7B-Instruct
FullContext 29.3 44.2 52.6 58.0 46.9 30.5 31.7 23.4 24.0 72.5 89.3 45.8 8.0 100.0 60.6 66.9 49.0
2K SnapKV 28.4 44.5 51.0 55.1 42.1 32.7 30.0 23.5 25.1 60.0 88.9 40.6 10.5 100.0 5.2 7.1 40.3
StreamingLLM 23.2 37.7 33.2 41.6 35.1 19.6 28.5 20.8 25.0 31.0 71.0 38.6 9.5 34.5 12.8 14.8 29.8
RAG 19.8 41.7 47.7 50.9 41.1 24.8 28.8 22.3 23.5 68.0 78.7 42.4 3.0 61.5 53.9 42.1 40.6
ILRe w/ $l_{R}=5$ 18.6 38.3 52.1 54.7 41.7 30.0 29.1 22.9 23.8 68.5 89.0 42.8 3.5 60.5 61.2 62.1 43.7
ILRe w/ $l_{R}=6$ 19.1 40.3 44.3 51.4 41.6 27.6 29.6 21.6 23.6 67.5 89.2 43.8 6.0 60.0 59.9 62.6 43.0
ILRe w/ $l_{R}=9$ 20.5 40.0 50.5 56.7 45.7 32.6 29.9 21.8 23.6 69.5 90.5 43.4 6.0 69.0 59.5 62.2 45.1
[…]
1K SnapKV 27.9 42.3 49.4 52.7 41.4 29.7 27.4 23.1 24.0 58.5 88.7 40.4 10.5 100.0 8.0 7.7 39.5
StreamingLLM 21.4 30.0 28.4 38.6 32.3 15.4 26.5 20.1 24.5 30.0 57.6 38.9 9.5 25.0 16.0 16.4 26.9
RAG 17.3 34.0 44.0 45.2 37.3 25.4 27.1 22.1 22.9 61.5 62.0 39.2 5.0 45.0 45.5 39.2 35.8
ILRe w/ $l_{R}=5$ 17.0 32.5 49.4 53.4 35.3 24.2 27.9 21.0 23.1 62.5 83.3 43.1 4.0 38.5 61.8 59.9 39.8
ILRe w/ $l_{R}=6$ 13.6 31.1 39.0 47.4 39.1 19.9 28.0 19.9 23.2 63.5 88.5 42.5 5.5 36.0 61.4 58.4 38.6
ILRe w/ $l_{R}=9$ 18.1 33.8 49.3 51.6 37.2 26.4 28.9 20.6 23.1 67.5 85.9 43.0 5.0 40.5 60.2 57.3 40.5
[…]
@@
We highlight two key observations. First, compared with other compression algorithms, ILRe demonstrates competitive performance at 2K token budget. ILRe with full KV retention attains the best overall performance, outperforming SnapKV and StreamingLLM in most tasks, with particularly strong gains in retrieval-heavy tasks such as QA and summarization. Even with various $l_{R}$, ILRe remains competitive, demonstrating that intermediate layer retrieval can recover useful long-context information with fewer stored tokens. Second, at 1K context length, ILRe continues to scale favorably, showing robustness under tighter memory budgets.
@@ 4.3 Ablations
Table 4: Different prefilling methods with Llama-3.1-UltraLong-8B-1M-Instruct in RULER.
@@
Prefilling Method. To demonstrate the stability of the streaming chunked prefill in our framework, we compare it with using the full KV in layers $\{1,2,...,l_{R}-1\}$ by setting $W=\infty$. As shown in Table 4, the results of RULER benchmark with the streaming chunked prefill are comparable to those of full KV. The results with “FullKV” in Table 3 further reinforce this finding.
Table 5: RULER benchmark performance of different pooling settings with Llama-3.1-UltraLong-8B-1M-Instruct. The best average scores are in bold faces.
@@
Multi Kernels. We define three different configurations of the pooling method from the default setups. Less-Max-Kernels uses only one max pooling with kernel size $4$. Less-Avg-Kernels reduces the number of average pooling to nine with sizes ranging from $1$ to $9$. One-Max-Avg only has one max pooling and one average pooling with size $4$ and $5$ respectively.
@@
Table 6: ILRe with/without Dual Chunk Attention in RULER benchmark. Because the long context measurements are time consuming, we halt the measurements after the score is below $30$. Qwen3-32B (Yang et al., 2025a) is measured only in the non-thinking mode.
Model Claimed Length Method 4K 8K 16K 32K 64K 128K 256K 512K 1M
Llama3.1-8B-Instruct 128K FullContext 95.7 93.9 91.4 86.6 83.8 74.8 OOM OOM OOM
ILRe w/ $l_{R}=3$ - 93.5 90.6 87.1 85.3 76.8 45.6 17.2 -
ILRe w/ $l_{R}=3$, DCA - 92.8 90.1 87.8 84.8 85.6 81.9 83.7 83.2
Qwen2.5-7B-Instruct 32K FullContext 94.5 92.6 91.8 89.0 69.4 25.8 OOM OOM OOM
ILRe w/ $l_{R}=5$ - 92.4 91.9 89.1 84.8 83.2 53.8 13.1 -
ILRe w/ $l_{R}=5$, DCA - 92.6 92.2 89.5 85.6 84.7 83.8 84.7 83.5
Qwen3-32B 32K FullContext 95.5 94.8 94.8 93.7 OOM OOM OOM OOM OOM
ILRe w/ $l_{R}=6$ - 94.7 93.0 89.7 86.7 85.0 20.4 - -
ILRe w/ $l_{R}=6$, DCA - 95.4 94.0 89.6 87.0 86.7 85.1 84.2 20.7
Length Extrapolation. We evaluate LLMs without lengthy tuning using the RULER benchmark (see Table 6), aiming to verify the compatibility of DCA with ILRe.
The default ILRe (without any extrapolation technique) performs well under $128K$ tokens across all tested models—even for the Qwen family with a native context window capacity of only $32K$ tokens. When DCA is enabled, ILRe further extends the effective context length of LLMs. Notably, results for the Llama family (presented in Table 1 and  6) demonstrate that Llama3.1-8B-Instruct, with both ILRe and DCA enabled, achieves results comparable to Llama-3.1-UltraLong-8B-1M-Instruct across the 128K1M128K–1M token range.